What is C if X = (A + B – C) / C and X = 0.2145, A = 142, and B = 6,658?
Answer #1

X = ( A + B - C ) / C

0.2145 = ( 142 + 6,658 - C ) / C

0.2145 C = 6,800 - C

1.2145 C = 6,800

C = 5,599.01 Approximately

Feel free to ask in case of any query relating to this question